OpenStudy (anonymous):

if you would write this on paper how would you write it 3n+16+7n

OpenStudy (anonymous):

3n + 16 =7n 3n+16=7n 3n+7n=16

OpenStudy (photon336):

You would combine like terms. 3n and 7n both have n's in them, the same base so you can add them both up. 3n+7n+16 10n+16

OpenStudy (photon336):

@bellasarah